# Attention-Based Multi-Scale Temporal Fusion Network for Uncertain-Mode Fault Diagnosis in Multimode Processes

### v1

@@ 3.4 Model implementation and application process
where $l=1,2,...,L$, $\bm{o}_{k}$ denotes the output of the fully connected layer, $F_{c}$ represents the fully connected layer. $c_{k,l}$ represents the $l$-th dimension of the output for the $k$-th sample, which indicates the probability that the sample belongs to the $l$-th category. In addition, dropout is added before the fully connected layer to prevent overfitting. The cross-entropy loss function is used as the objective function for the fault diagnosis model, and it is formulated as,
@@
where $N_{b}$ denotes the number of samples in a batch, and $y_{k,l}$ denotes the $l$-th dimension of the one-hot encoding of the true class label for the $k$-th sample.
The process of applying the AMTFN model to fault diagnosis is illustrated in Fig. 6. It mainly consists of two steps: the first is data collection and pre-processing, and the second is model construction, optimization, and evaluation. The details are as follows:
(1) Data collection and pre-processing mainly includes data acquisition, labeling, standardization, sliding window processing, and data splitting. Firstly, the monitoring data across multiple modes are collected, and each sample is assigned a corresponding category label. Then, z-score normalization is applied to standardize samples across all modes. The standardized result is formulated as $(\bm{x}_{k}^{\text{o}}-\bm{\mu})/\bm{\sigma}$, where $\bm{\mu}$ and $\bm{\sigma}$ denote the mean and standard deviation of the normal samples across all modes. And sliding window processing is employed to preserve dynamic temporal information. Finally, the data are divided into training, validation, and testing sets.
(2) Model construction and optimization involves training the fault diagnosis model using the training set and selecting the optimal model based on the validation set. Model evaluation is performed on the test set to assess the overall performance of the fault diagnosis model.
Fig. 6.: The AMTFN model for uncertain-mode fault diagnosis.
@@ 4.1 Evaluation metrics
Micro-F1, Macro-F1, Fault Diagnosis Rate(FDR), and False Positive Rate(FPR) are selected to evaluate the performance of the fault diagnosis model. To facilitate the presentation of the calculation formula for each evaluation metric, the confusion matrix for the $l$-th class is provided in Table 2, where each element represents the number of corresponding samples.
Table 2: Confusion matrix for the $l$-th class.
Predicted class is $l$
Predicted class is not $l$
Actual class is $l$ $TP_{l}$ $FN_{l}$
Actual class is not $l$ $FP_{l}$ $FN_{l}$
The Micro-F1 is computed as follows,
@@
The Macro-F1 is computed as follows,
@@
The FDR and FPR are computed as follows,
[…]
@@
FDR, also known as Recall, measures the proportion of samples with the true label $l$ that are correctly classified as $l$. FPR measures the proportion of samples whose true label is not $l$ that are misclassified as $l$.
In addition, to evaluate the scalability of the fault diagnosis model, the number of parameters, training time, and testing time for each model are also used for comparison.
@@ 4.2 Implementation details
The detailed structural parameters of each module in AMTFN is shown in Table 3. In the experiments, the learning rate is set to 0.01. The epoch and batch size are set to 30 and 512, respectively.
Table 3: Structure of the modules in AMTFN.
@@
To validate the effectiveness of the proposed AMTFN, several advanced fault diagnosis models are applied for comparison. The models used for comparison are as follows:
(1) DCNN [16]. The DCNN model is constructed using convolutional layers, pooling layers, dropout, and fully connected layers, aiming to capture spatiotemporal features.
(2) CNN-LSTM [18]. In the CNN-LSTM model, the CNN layer is used for feature learning, while the LSTM layer is responsible for capturing time delay information.
(3) IPO-ViT [19]. In the IPO-ViT model, the Transformer architecture is employed for global feature learning of complex process signals.
(4) Multi-scale generative and adversarial Metric networks (MGAMN) [34]. MGAMN was developed for cross-domain fault diagnosis. In this model, adversarial training and metric learning strategy is designed to learn common feature representations.
Methods 1-3 have demonstrated superior performance in single-mode fault diagnosis, while Method 4 represents a state-of-the-art domain generalization technique. Advanced single-mode and domain generalization fault diagnosis methods are applied to uncertain mode fault diagnosis scenario to validate the challenges in such settings. The comprehensive comparison of all methods is conducted to evaluate the performance of the proposed model.
The model settings for the ablation experiments are shown in Table 4.
Table 4: Model settings for the ablation experiments.
@@ 4.3 Case one: TE process
4.3.1 Dataset description and fault diagnosis tasks
The TE process simulation model was developed based on a real chemical process, which has become a benchmark for fault diagnosis and process monitoring [41]. It mainly consists of five components: the reactor, condenser, vapor-liquid separator, recycle compressor and product stripper. Bathelt et al. [42] introduced additional process measurements and disturbances to the TE process, which were implemented using SIMULINK. The overall structure of the extended TE process is shown in Fig. 7. The TE process includes 41 measured variables and 12 manipulated variables. It is capable of simulating 28 faults under 6 modes. Liu et al. obtained the multimode fault diagnosis datasets by adjusting the parameters of the TE process simulation models [43]. Since fault 3 and fault 16 are analogous with the normal operating conditions [44], they are not included in this study. The system health condition categories, modes and task settings are presented in Table 5 and Table 6. In tasks T1 to T6, each task’s dataset consists of data from five modes, with the corresponding positions marked as $\checkmark$.
@@ 4.3.1 Dataset description and fault diagnosis tasks
Table 5: Faults of TE process used in uncertain-mode fault diagnosis [45].
No. Description
N Normal
F1 A/C feed ratio, B composition constant (stream 4)
F2 B composition, A/C ratio constant (Stream 4)
F4 Reactor cooling water inlet temperature
F5 Condenser cooling water inlet temperature
F6 A feed loss (stream 1)
F7 C header pressure loss - reduced availability (stream 4)
F8 A, B, C feed composition (stream 4)
F9 D feed temperature (stream 2)
F10 C feed temperature (stream 4)
F11 Reactor cooling water inlet temperature
F12 Condenser cooling water inlet temperature
F13 Reaction kinetics
F14 Reactor cooling water valve
F15 Condenser cooling water valve
F17-20 Unknown
F21 The valve fixed at steady state position
Table 6: Modes of TE process used in uncertain-mode fault diagnosis.
@@
The TE dataset used in the experiment was generated through SIMULINK simulation of the TE process, covering 19 system health state categories across 6 modes. The simulation time was set to 100 hours, the sampling interval was set to 3 minutes and and the fault injection time was set to 30 hours. Note that during fault simulations in some modes, the simulation process might stop abruptly, resulting in fewer samples for some categories, leading to class imbalance. In sliding window processing, the window size $w$ was set to 64. For the data from each mode, a hierarchical splitting method was applied to divide the samples into training, validation, and testing sets in an 8:1:1 ratio.
4.3.2 Experimental results and analysis
The Micro F1 and Macro F1 scores of different methods on the TE process dataset are shown in Table 7 and Table 8, respectively. The proposed AMTFN achieves the highest average Micro F1 and average Macro F1 scores among the five fault diagnosis models, with values of 0.9792 and 0.9803, respectively. The DCNN has the lowest average Micro F1 and average Macro F1 scores, indicating that relying solely on convolution operations struggles to capture the complex features of multimode data distributions. Compared to the CNN-LSTM, IPO-ViT, and MGAMN models, AMTFN achieves a 3.18%, 9.08%, and 14.19% improvement in average Micro F1, respectively, and a 2.98%, 8.48%, and 15.47% improvement in average Macro F1, respectively. Taking Task T1 as an example, the two models with the highest average Micro F1 scores among the comparative models are selected for further comparison. The FDR and FPR scores of different methods on the TE process dataset are shown in Table 9. The proposed AMTFN model achieved the highest FDR, with FDR values exceeding 0.9 for both the N and F21 categories, whereas the comparative models showed FDR values below 0.8 for these categories. In addition, the AMTFN model achieved the lowest FPR. This validates that the proposed AMTFN model can effectively capture deep features in multimode distributed data, enabling efficient fault diagnosis under uncertain modes.
@@ 4.3.2 Experimental results and analysis
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
@@
T-distributed stochastic neighbor embedding (t-SNE) is used to visualize the feature distributions learned by each model. Taking task T1 as an example, the visualization results are shown in Fig. 8. From Fig. 8(a) and (d), it can be seen that the features learned by CNN and MGAMN reveal poorly separated clusters. Fault 14 is selected to compare CNN-LSTM, IPO-ViT, and the proposed model. The samples belonging to Fault 14 are highlighted with dashed boxes in Fig. 8(b), (c), and (e). It is evident that the features extracted by CNN-LSTM and IPO-ViT are too scattered, whereas AMTFN effectively clusters the samples belonging to this category. This demonstrates that the features extracted by AMTFN exhibit minimal differences across modes, highlighting its superior cross-mode performance.
Fig. 8.: Feature map visualization results via t-SNE on task T1.
To evaluate the scalability of different models, the number of parameters, training time per epoch, and testing time per epoch for each model are presented in Table 10. AMTFN has the smallest number of parameters and the shortest training time, with testing time exceeding that of MGACN by only 0.28 seconds. This is because the GRU in AMTFN limits parallel processing, which adds the computation time. However, AMTFN demonstrates unparalleled advantages in inferring system health conditions. The computational complexity of DCNN, CNN-LSTM, and IPO-ViT is significantly higher than that of AMTFN, which highlights the advantages of AMTFN in practical applications.
Table 10: Parameter number, training time and test time of different methods on the TE process dataset.
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
@@ 4.3.3 Ablation study
To verify the effectiveness of the MSDC, GRU, and TAM, ablation experiments were conducted. The Micro F1 scores of different methods are shown in Table 11. The average Micro F1 score of A2 is 17.56% higher than that of A1, indicating that the GRU is superior to the CNN in temporal feature extraction. The model A3 and A4 are obtained by adding TAM to A1 and A2, respectively (see Table 4). The average Micro F1 score of A3 is 10.19% higher than that of A1, and A4 is 1.63% higher than that of A2. This demonstrates that TAM effectively fuses features extracted by both CNN and GRU. A5 combines A1 and A2, meaning that both the MSDC and GRU are used for feature extraction. The average Micro F1 score of A5 is 20.35% higher than that of A1 and 2.38% higher than that of A2. This demonstrates that combining MSDC and GRU enhances the ability to capture the complex structures within the data. AMTFN is equivalent to adding GRU to A3, or adding MSDC to A4, or adding TAM to A5. The average Micro F1 score of TAGRUN is 9.35% higher than A3, 0.84% higher than A4, and 0.11% higher than A5. This indicates that GRU, MSDC, and TAM complement each other, collectively improving the model’s performance in uncertain-mode fault diagnosis.
Since the improvement in AMTFN’s average Micro F1 score over A4 and A5 is relatively small, the feature maps extracted from A4 and A5 are visualized using t-SNE. Taking Task T1 as an example, the visualization results are shown in Fig. 9. Similarly, the samples belonging to Fault 14 are highlighted with dashed circles. Although A4 achieves a Micro F1 score of 0.9707, the features of samples from different categories exhibit poor separability. A5 effectively clusters the features of samples from different categories into different groups. However, the features of samples within the same category exhibit significant dispersion. This may be because A5 learns mode-specific features, causing each operation mode and system health condition category pair to be clustered into a unique group as tightly as possible. For uncertain-mode fault diagnosis, it is essential to ensure both the separability of samples from different categories and the compactness of samples from the same category. The proposed AMTFN model achieves better intra-category compactness and maintains a reasonable degree of inter-category separability. This demonstrates that focusing on critical time points through TAM may help the model capture invariant features across different modes, thereby reducing feature differences among samples of the same category.
@@
Fig. 9.: Feature map visualization results via t-SNE on task T1.
@@ 4.4 Case two: three-phase flow facility
4.4.1 Dataset description and fault diagnosis tasks
In the field of fault diagnosis, the dataset from three-phase flow facility [46] at Cranfield University has been used extensively as a benchmark for evaluating the performance of various methods. The three-phase flow facility is designed to provide controlled flows of water, oil, and air. Its structure is shown in Fig. 10. This facility captured 24 process variables and simulated six faults. The operation mode settings and the dataset used in the experiments are presented in Table 12, where different water flow rates and air flow rates correspond to different operating modes. The partitioning ratios of the training, validation, and test sets, as well as the sliding window settings, are consistent with those set in the TE process.
Fig. 10.: Sketch of the three-phase flow facility. [46].
Table 12: Faults of three-phase flow facility used in uncertain-mode fault diagnosis.
@@ 4.4.1 Dataset description and fault diagnosis tasks
4.4.2 Experimental results and analysis
The FDR, FPR, Micro F1 and Macro F1 scores of different methods on the three-phase flow facility dataset are shown in Table 13. The proposed AMTFN achieves the highest FDR, average Micro F1, and average Macro F1 scores, all with values of 1.0. In addition, the AMTFN model achieved the lowest FPR, with a value of 0. Compared to DCNN, CNN-LSTM, IPO-ViT, and MGAMN, AMTFN’s average FDR is higher by 25.15%, 0.01%, 6.93%, and 8.51%, respectively; its average Micro F1 is higher by 12.21%, 0.03%, 3.75%, and 4.68%, respectively; and its Macro F1 is higher by 35.10%, 0.03%, 6.18%, and 7.68%, respectively. For the system health condition category N, the FDR scores of DCNN, IPO-ViT, and MGAMN are all below 0.72, whereas both AMTFN and CNN-LSTM achieve a score of 1. Additionally, CNN-LSTM achieves an FDR of 0.9993 for category F3, whereas AMTFN still achieves a score of 1. Overall, the AMTFN model outperforms the other models across all four metrics, further validating its superiority in uncertain-mode fault diagnosis.
Table 13: FDR, FPR, Micro F1 and Macro F1 scores of different methods on the three-phase flow facility dataset.

### v2

@@ 3.4 Model output and application process
where $l=1,2,...,L$, $F_{c}$ denotes the fully connected layer. $c_{k,l}$ represents the $l$-th output component of the output for the $k$-th sample, which indicates the predicted confidence that the sample is assigned to category $l$. In addition, dropout is added to prevent overfitting. The cross-entropy loss is applied as the objective function for the fault diagnosis model, and it is defined as,
@@
where $N_{b}$ denotes the batch size, and $y_{k,l}$ takes the value 0 or 1, which indicates whether the label of the $k$-th sample is $l$.
The process of applying the AMTFNet model to fault diagnosis is illustrated in Fig. 6, which mainly includes two steps: the first is data acquisition and preliminary processing, and the second is model construction, optimization, and evaluation. The detailed description is provided below.
(1) Data acquisition and preliminary processing mainly includes data acquisition, labeling, standardization, sliding window processing, and data splitting. Firstly, the monitoring data across multiple modes are collected, and each sample is assigned a corresponding category label. Then, z-score normalization is applied to standardize samples across all modes. The standardized result is formulated as $(\bm{x}_{k}^{\text{o}}-\bm{\mu})/\bm{\sigma}$, where $\bm{\mu}$ and $\bm{\sigma}$ denote the mean and standard deviation estimated from the fault-free samples across all modes. And sliding window of size 64 is employed to preserve dynamic temporal information. Finally, the data are divided into training, validation, and test sets at a ratio of 8:1:1.
(2) Model construction and optimization involves training the fault diagnosis model and selecting the best-performing model. Model evaluation is performed on the test set to measure the overall effectiveness of the fault diagnosis model.
Fig. 6.: The process of applying AMTFNet model in uncertain-mode fault diagnosis.
@@ 4.1 Evaluation metrics
Micro-F1, Macro-F1, Fault Diagnosis Rate (FDR), and False Positive Rate (FPR) are selected to compare the model’s ability in diagnosing faults. To facilitate the presentation of the calculation formula for each evaluation metric, the confusion matrix corresponding to the $l$-th category is provided in Table 2, where each element denotes the quantity of data instances matching the relevant condition.
Table 2: Confusion matrix corresponding to the $l$-th category.
Inferred category is $l$
Inferred category is not $l$
Real category is $l$ $TP_{l}$ $FN_{l}$
Real category is not $l$ $FP_{l}$ $FN_{l}$
The Micro-F1 is computed as follows [42],
@@
The Macro-F1 is computed as follows [42],
@@
The FDR and FPR are computed as follows [17],
$\text{FDR}_{l}$ $=\frac{\mathrm{TP}_{l}}{\mathrm{TP}_{l}+\mathrm{FN}_{l}}$ (11)
@@
FDR measures the ratio of instances with the real category $l$ that are correctly classified as $l$. FPR measures the ratio of instances from other categories are mistakenly classified as category $l$.
Additionally, both model size and runtime are considered as metrics to examine the scalability of the fault diagnosis methods.
@@ 4.2 Implementation details
The detailed structure of the modules in AMTFNet is shown in Table 3. The training procedure uses a learning rate of 0.01, 30 epochs and a batch size of 512.
Table 3: Structure of the modules in AMTFNet.
@@
To demonstrate the capability of the proposed AMTFNet, several advanced models are applied for comparison. The models used for comparison are as follows:
(1) DCNN [17]. The CNN is utilized in this model to capture deep representative features.
(2) CNN-LSTM [19]. CNN and LSTM are fused in this model to collaboratively capture features indicative of health conditions.
(3) IPO-ViT [20]. This model employed the Transformer-based structure to capture global features.
(4) MGAMN [36]. This model was constructed by combining data augmentation and representation learning for fault diagnosis under domain shift conditions.
Methods 1-3 have demonstrated superior performance in SMFD, while Method 4 represents a state-of-the-art DGFD. Advanced SMFD and DGFD methods are employed in UMFD scenario to validate the challenges in such settings. The comprehensive comparison of all methods is conducted to validate the effectiveness of the developed model.
The setup details for the ablation experiments are shown in Table 4.
Table 4: Setup details for the ablation experiments.
@@ 4.3 Case one: TE process
4.3.1 Task description
The TE process simulation model has become a benchmark for industrial system fault diagnosis [43]. Bathelt et al. [44] introduced additional process measurements and disturbances to the TE process, as shown in Fig. 7. This process includes 41 measured variables and 12 manipulated variables. It is capable of simulating 28 faults under 6 modes. Liu et al. obtained the multimode fault diagnosis datasets by adjusting the parameters of the TE process simulation models [45]. Since fault 3 and fault 16 are analogous with the normal operating conditions [46], they are not included in this study. The system health condition categories, modes and task settings are presented in Table 5 and Table 6. In tasks T1 to T6, each task’s dataset consists of data from five modes, with the corresponding positions marked as $\checkmark$.
@@ 4.3.1 Task description
Table 5: Faults of TE process used in uncertain-mode fault diagnosis [47, 44].
No. Description Type
N Normal -
F1 A/C feed ratio, B composition constant (stream 4) Step
F2 B composition, A/C ratio constant (Stream 4) Step
F4 Reactor cooling water inlet temperature Step
F5 Condenser cooling water inlet temperature Step
F6 A feed loss (stream 1) Step
F7 C header pressure loss - reduced availability (stream 4) Step
F8 A, B, C feed composition (stream 4) Random variation
F9 D feed temperature (stream 2) Random variation
F10 C feed temperature (stream 4) Random variation
F11 Reactor cooling water inlet temperature Random variation
F12 Condenser cooling water inlet temperature Random variation
F13 Reaction kinetics Drift
F14 Reactor cooling water valve Sticking
F15 Condenser cooling water valve Sticking
F17-20 Unknown Unknown
F21 A feed temperature (stream1) Random variation
Table 6: Modes of TE process used in uncertain-mode fault diagnosis [47].
@@
The TE dataset used in the experiment was generated through SIMULINK simulation of the TE process, covering 19 system health state categories across 6 modes. The simulation was run for 100 hours with samples collected every 3 minutes and the fault was introduced at the 30th hour. Note that during fault simulations in some modes, the simulation process might stop abruptly, resulting in fewer samples for some categories, leading to class imbalance.
4.3.2 Experiment results
The experiment results for Micro F1 and Macro F1 scores are presented in Table 7 and Table 8, respectively. AMTFNet achieves the highest average Micro F1 and average Macro F1 scores among the five fault diagnosis models, with values of 0.9792 and 0.9803, respectively. The DCNN has the lowest average Micro F1 and average Macro F1 scores, indicating that adopting only convolution operations struggles to capture the complex features of multimode data distributions. Compared to the CNN-LSTM, IPO-ViT, and MGAMN models, AMTFNet achieves a 3.18%, 9.08%, and 14.19% increase in average Micro F1 scores, respectively, and a 2.98%, 8.48%, and 15.47% improvement in average Macro F1 scores, respectively. Taking Task T1 as an illustration, the two models with the highest average Micro F1 scores among the comparative models are selected for further comparison. The experiment results for FDR and FPR scores are shown in Table 9. AMTFNet model achieved the highest FDR, with FDR values exceeding 0.9 for both the N and F21 categories, whereas the comparative models showed FDR values below 0.8 for these categories. In addition, the AMTFNet model achieved the lowest FPR. This validates that the proposed AMTFNet model can effectively capture deep features in multimode distributed data, enabling efficient fault diagnosis under uncertain modes.
@@ 4.3.2 Experiment results
DCNN CNN-LSTM IPO-ViT MGAMN AMTFNet
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFNet
@@
The feature distributions learned by each model are visualized through t-SNE. Taking task T1 as an example, the details of the visual representation are presented in Fig. 8. Fig. 8(a) and (d) reveal that the features learned by CNN and MGAMN reveal poorly separated clusters. Fault 14 is selected to compare CNN-LSTM, IPO-ViT, and the proposed model. The samples belonging to Fault 14 are highlighted with dashed boxes in Fig. 8(b), (c), and (e). It is evident that the features extracted by CNN-LSTM and IPO-ViT are too scattered, whereas AMTFNet effectively clusters the samples belonging to this category. This demonstrates that the features extracted by AMTFNet exhibit minimal differences across modes, highlighting its superior cross-mode performance.
Fig. 8.: T-SNE visualization results for task T1 on the TE process dataset.
The experiment results for scalability comparison are presented in Table 10. AMTFNet has the smallest number of parameters and the shortest training time, with testing time exceeding that of MGACN by only 0.28 seconds. This is because the GRU in AMTFNet limits parallel processing, which adds the computation time. However, AMTFNet demonstrates unparalleled advantages in inferring system health conditions. The computational complexity of DCNN, CNN-LSTM, and IPO-ViT is significantly higher than that of AMTFNet, which highlights the advantages of AMTFNet in practical applications.
Table 10: Experiment results for scalability comparison on the TE process dataset.
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFNet
@@ 4.3.3 Ablation study
To examine the capability of the MSDC, GRU, and TAM, ablation experiments were conducted. The Micro F1 scores for different ablation models are presented in Table 11. The average Micro F1 score of A2 is 17.56% higher than that of A1, indicating that the GRU is superior to the CNN in temporal feature extraction. The model A3 and A4 are obtained by adding TAM to A1 and A2, respectively (see Table 4). The average Micro F1 score of A3 is 10.19% higher than that of A1, and A4 is 1.63% higher than that of A2. This demonstrates that TAM effectively fuses features extracted by both CNN and GRU. A5 combines A1 and A2, meaning that both the MSDC and GRU are jointly employed to extract features. The average Micro F1 score of A5 is 20.35% higher than that of A1 and 2.38% higher than that of A2. This demonstrates that combining MSDC and GRU enhances the power to learn the complex features within the data. AMTFNet is equivalent to adding GRU to A3, or adding MSDC to A4, or adding TAM to A5. The average Micro F1 score of TAGRUN is 9.35% higher than A3, 0.84% higher than A4, and 0.11% higher than A5. This indicates that GRU, MSDC, and TAM complement each other, collectively strengthening the model’s capability for uncertain-mode fault diagnosis.
Since the improvement in AMTFNet’s average Micro F1 score over A4 and A5 is relatively small, the feature maps extracted from A4 and A5 are visualized using t-SNE. Taking Task T1 as an example, the visualization results are shown in Fig. 9. Similarly, the samples belonging to Fault 14 are highlighted with dashed circles. Although A4 achieves a Micro F1 score of 0.9707, the feature distributions of various classes exhibit poor separability. A5 effectively clusters the feature distributions of various classes into different groups. However, the features of samples within the same category exhibit significant dispersion. This may be because A5 learns mode-specific features, causing each operation mode and system health condition category pair to be clustered into a unique group as tightly as possible. For uncertain-mode fault diagnosis, it is essential to ensure both the separability of various categories and the compactness of the same category. The proposed AMTFNet model achieves better intra-category compactness and maintains a reasonable degree of inter-category separability. This demonstrates that focusing on critical time points through TAM may help the model capture invariant features across different modes, thereby reducing feature differences among samples of the same category.
Table 11: Micro F1 scores of different ablation models on the TE process dataset.
@@
Fig. 9.: T-SNE visualization results for ablation study on the TE process dataset.
@@ 4.4 Case two: three-phase flow facility
4.4.1 Task description
In the field of fault diagnosis, the dataset from three-phase flow facility (TPFF) [48] at Cranfield University has been commonly adopted to assess the effectiveness of various methods. The three-phase flow facility is designed to provide controlled flows of water, oil, and air. Its structure is shown in Fig. 10. This facility captured 24 process variables and simulated six faults. The operation mode settings and the dataset used in the experiments are presented in Table 12, where different water flow rates and air flow rates correspond to different operating modes.
Fig. 10.: Sketch of the TPFF. [48].
Table 12: Faults of TPFF used in uncertain-mode fault diagnosis [48].
@@ 4.4.1 Task description
4.4.2 Experiment results
The experiments for FDR, FPR, Micro F1 and Macro F1 scores are shown in Table 13. AMTFNet obtains the highest FDR, average Micro F1, and average Macro F1 scores, all with values of 1.0. In addition, the AMTFNet model achieved the lowest FPR, with a value of 0. Compared to DCNN, CNN-LSTM, IPO-ViT, and MGAMN, AMTFNet’s average FDR is higher by 25.15%, 0.01%, 6.93%, and 8.51%, respectively; its average Micro F1 is higher by 12.21%, 0.03%, 3.75%, and 4.68%, respectively; and its Macro F1 is higher by 35.10%, 0.03%, 6.18%, and 7.68%, respectively. For the system health condition category N, the FDR scores of DCNN, IPO-ViT, and MGAMN are all below 0.72, whereas both AMTFNet and CNN-LSTM achieve a score of 1. Additionally, CNN-LSTM achieves an FDR of 0.9993 for category F3, whereas AMTFNet still achieves a score of 1. Overall, the AMTFNet model outperforms the other models across all four metrics, further validating its superiority in uncertain-mode fault diagnosis.
Table 13: FDR, FPR, Micro F1 and Macro F1 scores of different methods on the TPFF dataset.